# Are Images Indistinguishable to Humans Also Indistinguishable to Classifiers?

### v3

@@ Abstract
The ultimate goal of generative models is to perfectly capture the data distribution. For image generation, common metrics of visual quality (e.g., FID) and the perceived truthfulness of generated images seem to suggest that we are nearing this goal. However, through distribution classification tasks, we reveal that, from the perspective of neural network-based classifiers, even advanced diffusion models are still far from this goal. Specifically, classifiers are able to consistently and effortlessly distinguish real images from generated ones across various settings. Moreover, we uncover an intriguing discrepancy: classifiers can easily differentiate between diffusion models with comparable performance (e.g., U-ViT-H vs. DiT-XL), but struggle to distinguish between models within the same family but of different scales (e.g., EDM2-XS vs. EDM2-XXL). Our methodology carries several important implications. First, it naturally serves as a diagnostic tool for diffusion models by analyzing specific features of generated data. Second, it sheds light on the model autophagy disorder and offers insights into the use of generated data: augmenting real data with generated data is more effective than replacing it.
@@ 1 Introduction
Diffusion probabilistic models (Sohl-Dickstein et al., 2015; Song et al., 2020; Ho et al., 2020) have emerged as leading generative models for image (Ramesh et al., 2021; Esser et al., 2024), video (Brooks et al., 2024; Bao et al., 2024; Zhao et al., 2024) and 3D content generation (Poole et al., 2022), surpassing previous generative models such as generative adversarial networks (Goodfellow et al., 2014), variational autoencoders (Kingma & Welling, 2013), and normalizing flows (Rezende & Mohamed, 2015). In particular, focusing on image generation, diffusion models are capable of producing realistic images that are often indistinguishable from real ones to the human eye (see Fig. 2). Furthermore, common metrics such as FID (Heusel et al., 2017) suggest that the distribution generated by state-of-the-art diffusion models closely approximates the ImageNet validation set (Karras et al., 2023; 2022; Tian et al., 2024). Therefore some works, such as Wang et al. (2024); You et al. (2024), have utilized generated data to augment training datasets. However, to the best of our knowledge, few studies have been conducted to analyze the differences between the generated and real distributions.
This raises a fundamental question: How far is the generated distribution from the real distribution? Inspired by the methodology in the work (Liu & He, 2024; Tommasi et al., 2017) on dataset bias (Torralba & Efros, 2011), we propose distribution classification tasks (see Sec. 3 for details), predicting the distribution an image belongs to - real or generated - thereby serving as a probe to explore this critical question. Notably, binary classification is a natural approach, as classifiers parameterized by neural networks can serve as effective indicators of the differences between the real and generated distributions through classification accuracy (see more analysis in Appendix  D). Compared to the widely used FID metric, our classifier-based approach directly measures distribution differences without relying on Gaussian assumptions (see further discussion in Appendix E).
@@
Furthermore, we observe intriguing contradictions between classifiers’ prediction and both FID and human judgments through two groups of experiments. The first group involves classifying the smallest and largest models within the same diffusion family, which share inductive biases but differ in visual quality. The second group examines diffusion models with similar FID scores and visual quality but different inductive biases. Surprisingly, classifiers perform well on the second task but struggle with the first, revealing contradictions between classifier performance and FID scores. Meanwhile, human judgments align with classifiers in the first task but degrade to random guessing in the second, indicating inconsistency between classifier and human judgments.
Our approach displays several important implications. First, it naturally serves as a diagnostic tool for diffusion models by extending the analysis to specific features of generated data. For example, in evaluating the advanced diffusion transformer U-ViT-H/2 (Bao et al., 2023), we focus on two key aspects: spatial features and frequency ranges. Spatially, classifiers can easily distinguish whether an image is real or generated using any part of the image. In terms of frequency, classifiers struggle with minimal low-frequency information but perform well across all other frequency intervals. This suggests that U-ViT-H/2 has effectively learned low-frequency features, making generated images harder to distinguish in these regions, but remains less proficient in higher frequency regions. Second, our approach enables the evaluation of using generated data for training models. It provides a reasonable explanation for the model autophagy disorder (MAD) phenomenon observed in Shumailov et al. (2024); Alemohammad et al. (2024). This collapse is surprising because, despite the realistic appearance of the generated data, using it in a continuous loop to train subsequent generations of models leads to failure. We explain this by showing that there exists a distribution mismatch between the training and generated distributions in each generation. As this mismatch accumulates, the generated data distribution drifts further from the original. In downstream tasks, such as supervised and semi-supervised classification, this mismatch leads to performance degradation when real data is simply replaced by generated data. However, this distribution drift can be mitigated by augmenting real data with generated data, which introduces novel features while maintaining alignment with the original data distribution, leading to improved performance.
@@
  • We provide a reasonable explanation for the surprising model autophagy disorder phenomenon and show that augmenting real data with generated data is more effective than replacing it in supervised and semi-supervised learning.
2 Background
2.1 Diffusion model
Diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2020; Ho et al., 2020) gradually inject noise into data $\bm{x}$ during the forward process:
$\bm{z}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{x}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},$ (1)
and remove noise to generate data in the reverse process. Diffusion models typically use a noise prediction network $\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t)$ to predict the noise $\bm{\epsilon}$ added to $\bm{z}_{t}$. Noise prediction loss is defined as:
$\mathcal{L}=\mathbb{E}_{t,\bm{x}_{0},\bm{\epsilon}}[\|\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t)-\bm{\epsilon}\|_{2}^{2}].$ (2)
2.2 Frequency filter
The frequency content of an image represents the rate of pixel value changes. Low-frequency components capture overall shapes and gradual grayscale changes, while high-frequency components reflect fine details like edges and textures. Notably, convolutional neural networks can detect high-frequencies that are often imperceptible to humans (Wang et al., 2020a). In our paper, we implement a rectangular filter and discuss the choice between using rectangular or circular filters in Appendix F.
Low-pass filters are implemented as follows: Compute the Fourier transform of the image $f(x,y)$ using $F(u,v)=\text{FFT}(f(x,y))$, and then center the spectrum with $F_{c}(u,v)=\text{fftshift}(F(u,v))$. Define a rectangular mask $H(u,v)$ in the frequency domain, where $M$ and $N$ are the image dimensions:
$H(u,v)=\begin{cases}1&\text{if }\left|u-\frac{M}{2}\right|\leq\text{threshold}\text{ and }\left|v-\frac{N}{2}\right|\leq\text{threshold},\\ 0&\text{otherwise}.\end{cases}$ (3)
$H(u,v)=\begin{cases}0&\text{if }\left|u-\frac{M}{2}\right|\leq\text{threshold}\text{ and }\left|v-\frac{N}{2}\right|\leq\text{threshold},\\ 1&\text{otherwise}.\end{cases}$ (4)
@@ 3 Related work
Dataset bias. Torralba & Efros (2011) introduced the "Name That Dataset" game to discover dataset bias. Tommasi et al. (2017) and Liu & He (2024) expanded on this by applying convolutional neural networks and large-scale datasets, respectively. Existing work on dataset bias focuses on classification to highlight biases in data collection for building unbiased datasets, our work leverages distribution classification to explore a fundamental question: How far is the generated distribution from the real distribution? Besides, we focus on the discrepancy between the generated distribution and the real distribution rather than between specific datasets. Furthermore, the conclusion drawn by related work highlights the ability of neural networks to detect biases within increasingly general datasets, our conclusion underscores the significant gap that still exists between current diffusion models and the real data distribution. We also highlight the potential of distribution classification methods for gaining deeper insights into generative models and more effectively leveraging generated data.
Generated image detection. Several empirical works on generated image detection span three key areas: universal detection (Liu et al., 2024; Ojha et al., 2023; Wang et al., 2020b; Girish et al., 2021), reconstruction-based methods (Wang et al., 2023b; Luo et al., 2024), and frequency analysis (Wang et al., 2020a; Zhang et al., 2019; Durall et al., 2020; Chandrasegaran et al., 2021; Frank et al., 2020; Dzanic et al., 2020; Yang et al., 2023; Corvi et al., 2023b; Ricker et al., 2022; Corvi et al., 2023a). Existing work generally focuses on either building universal discriminators to detect generated images or identifying frequency artifacts to build classifiers. In contrast, our approach leverages distribution classification to gain deeper insights into generative models and quantify the discrepancy between generated and real distributions. Instead of prioritizing classifier performance, we focus on the insights provided into generative model behavior.
@@ 4 Experimental settings
We present the main experimental settings as follows. For more details, please see Appendix A.
Dataset. For label-to-image tasks, we use the CIFAR-10 (Krizhevsky et al., 2009) and ImageNet (Deng et al., 2009) datasets, and for text-to-image tasks, we utilize the COCO2014 dataset (Lin et al., 2014). For CIFAR-10, the real distribution is constructed from its training and testing sets, comprising 50k training images and 10k testing images. The generated distribution is created using diffusion models to produce an equivalent number of images. For ImageNet, we consider resolutions of 256 and 512 and follow ADM’s (Dhariwal & Nichol, 2021) method to preprocess ImageNet into the desired resolutions. We randomly sample 100k training images and 50k validation images from the corresponding ImageNet sets to represent the real distribution, and generate an equivalent number of images using diffusion models for the generated distribution. For COCO2014, we randomly sample 10k training images and 1k validation images from the respective sets to form the real distribution, and use the corresponding prompts to generate an equivalent number of data for the generated distribution.
Classifier. By default, we use ResNet-50 (He et al., 2016) as the classifier architecture, and also consider ConvNeXt-T (Liu et al., 2022) and ViT-S (Dosovitskiy et al., 2020) for completeness. Our pre-processing follows standard supervised training (Liu et al., 2022). Specifically, during training, the classifiers process randomly augmented crops of $224\times 224$ images. For validation, images are resized to 256 pixels on the shorter side, preserving the aspect ratio, and then center cropped to $224\times 224$. For CIFAR-10 experiments, we initialize the classifier with ResNet-50 pre-trained on ImageNet and also provide the results trained from scratch (see Appendix C.2). For ImageNet, we train ResNet-50 from scratch.
Diffusion model. For CIFAR-10 generation, we consider two diffusion models: EDM (Karras et al., 2022) and U-ViT (Bao et al., 2023). For ImageNet-256, we explore three diffusion models: EDM2 (Karras et al., 2023), U-ViT-H/2 (Bao et al., 2023), and DiT-XL/2 (Peebles & Xie, 2023). Since EDM2 is originally designed for ImageNet-512 generation, we resize its generated images from 512 to 256 resolution, with EDM2-XXL achieving an FID of 2.14, comparable to U-ViT-H/2 and DiT-XL/2. For ImageNet-512 generation, we use EDM2-XXL (Karras et al., 2023). For COCO generation, we consider three text-to-image diffusion models: Pixart-$\alpha$ (Chen et al., 2023), SDXL (Podell et al., 2023), and Playground-v2.5 (Li et al., 2024).
@@
Abbreviations. In the figures and tables following this paper, the following abbreviations are used: CIFAR-10 (C), ImageNet (I), U-ViT (U), EDM (E), DiT-XL/2 (D).
@@ 5 Generated distributions are easily classified as generated
In this section, we explore how the distribution generated by diffusion models, known for their success in image generation, differs from the real distribution. To analyze this, we use a distribution classification task to assess the extent of the difference.
Table 1: Binary distribution classification on the same distribution. Classifiers are indeed unable to distinguish samples from the same distribution, initially demonstrating the validity of our approach.
Distribution1 Distribution2 Classifier Accuracy (%)
U (Bao et al., 2023) U ResNet-50 50.83
ViT-S 50.57
ConvNeXt-T 50.18
E (Karras et al., 2022) E ResNet-50 50.67
ViT-S 50.44
ConvNeXt-T 50.14
@@ 5.2 Classifiers achieve high accuracy across various settings
Label-to-image. As shown in Tab. 2, we report the binary distribution classification accuracy for various combinations of real and generated distributions in the label-to-image scenario. Surprisingly, despite the low FID scores achieved by diffusion models, neural network-based classifiers still identify significant differences between the generated and real distributions. Specifically, all classifiers easily distinguish real from generated images across different dataset combinations, with most achieving over 98% accuracy in distribution classification. Moreover, we observe a positive correlation between distribution classification accuracy and the number of training samples for the combination of U-ViT-H/2 and ImageNet-256 (see Fig. 2). This suggests that as sample size increases, classifiers learn more generalizable features rather than merely memorizing the data. Notably, with only 50k samples per distribution (less than 10% of the ImageNet samples, approximately 1.28 million, used to train the generator), classifiers achieve over 99.5% accuracy. Besides, we extend our evaluation to other generative models, e.g., GANs (see detailed results in Sec. C.1).
Table 2: Binary distribution classification on label-to-image. All classifiers yield high accuracy on various combinations of datasets and generative models. FIDs are taken from related references.
Real dataset Generative model Corrresponding FID Classifier Accuracy (%)
C U (Bao et al., 2023) 3.11 ResNet-50 99.92
@@
Table 3: Four-way distribution classification on text-to-image. All classifiers yield high accuracy in distinguishing four distributions: COCO (Lin et al., 2014), Pixart-$\alpha$ (Chen et al., 2023), SDXL (Podell et al., 2023), and Playground-v2.5 (Li et al., 2024), using only 5k or 10k training samples per dataset. Notably, as the number of training samples increases, the accuracy of distribution classification consistently improves.

### v4

@@ Abstract
The ultimate goal of generative models is to perfectly capture the data distribution. For image generation, common metrics of visual quality (e.g., FID) and the perceived truthfulness of generated images seem to suggest that we are nearing this goal. However, through distribution classification tasks, we reveal that, from the perspective of neural network-based classifiers, even advanced diffusion models are still far from this goal. Specifically, classifiers are able to consistently and effortlessly distinguish real images from generated ones across various settings. Moreover, we uncover an intriguing discrepancy: classifiers can easily differentiate between diffusion models with comparable performance (e.g., U-ViT-H vs. DiT-XL), but struggle to distinguish between models within the same family but of different scales (e.g., EDM2-XS vs. EDM2-XXL). Our methodology carries several important implications. First, it naturally serves as a diagnostic tool for diffusion models by analyzing specific features of generated data. Second, it sheds light on the model autophagy disorder and offers insights into the use of generated data: augmenting real data with generated data is more effective than replacing it. Third, classifier guidance can significantly enhance the realism of generated images.
@@ 1 Introduction
Diffusion probabilistic models [52, 53, 26] have emerged as leading generative models for image [45, 17], video [5, 4, 69] and 3D content generation [44], surpassing previous generative models such as generative adversarial networks [21], variational autoencoders [29], and normalizing flows [46]. In particular, focusing on image generation, diffusion models are capable of producing realistic images that are often indistinguishable from real ones to the human eye (see Fig. 2). Furthermore, common metrics such as FID [25] suggest that the distribution generated by state-of-the-art diffusion models closely approximates the ImageNet validation set [28, 27, 55]. Therefore some works [62, 65], have utilized generated data to augment training datasets. However, to the best of our knowledge, few studies have been conducted to analyze the differences between the generated and real distributions.
This raises a fundamental question: How far is the generated distribution from the real distribution? Drawing inspiration from methodologies in dataset bias [34, 57, 58], we propose distribution classification tasks (see Sec. 2 for details) that predict whether an image belongs to the real or generated distribution. These tasks serve as an effective probe to quantify and analyze this distributional gap. We use the term “distribution classification” because generative models aim to capture an entire distribution rather than a finite dataset, allowing infinite sampling.
Notably, binary classification is a natural approach, as classifiers parameterized by neural networks can serve as effective indicators of the differences between the real and generated distributions through classification accuracy (see more analysis in Appendix  E). Compared to the widely used FID metric, our classifier-based approach directly measures distribution differences without relying on Gaussian assumptions (see further discussion in Appendix F).
@@
Furthermore, we observe intriguing contradictions between classifiers’ prediction and both FID and human judgments through two groups of experiments. The first group involves classifying the smallest and largest models within the same diffusion family, which share inductive biases but differ in visual quality. The second group examines diffusion models with similar FID and visual quality but different inductive biases. Surprisingly, classifiers perform well on the second task but struggle with the first, revealing contradictions between classifier performance and FID. Meanwhile, human judgments align with classifiers in the first task but degrade to random guessing in the second, indicating inconsistency between classifier and human judgments.
Our approach displays several important implications. First, it naturally serves as a diagnostic tool for diffusion models by extending the analysis to specific features of generated data. For example, in evaluating the advanced diffusion transformer U-ViT-H/2 [3], we focus on two key aspects: spatial features and frequency ranges. Spatially, classifiers can easily distinguish whether an image is real or generated using any part of the image. In terms of frequency, classifiers struggle with minimal low-frequency information but perform well across all other frequency intervals. This suggests that U-ViT-H/2 has effectively learned low-frequency features, making generated images harder to distinguish in these regions, but remains less proficient in higher-frequency regions. Second, our approach enables the evaluation of using generated data for training models. It provides a reasonable explanation for the model autophagy disorder (MAD) phenomenon observed in previous work [50, 1]. This collapse is surprising because, despite the realistic appearance of the generated data, using it in a continuous loop to train subsequent generations of models leads to failure. We explain this by showing that there exists a distribution mismatch between the training and generated distributions in each generation. As this mismatch accumulates, the generated data distribution drifts further from the original. In downstream tasks, such as supervised and semi-supervised classification, this mismatch leads to performance degradation when real data is simply replaced by generated data. However, this distribution drift can be mitigated by augmenting real data with generated data, which introduces novel features while maintaining alignment with the original data distribution, leading to improved performance.
@@
  • We provide a reasonable explanation for the model autophagy disorder phenomenon and show that augmenting real data with generated data is more effective than replacing it in supervised and semi-supervised learning.
  • We demonstrate that classifier guidance can be used to enhance the realism of generated images.
@@ 2 Related work
Dataset bias. Torralba and Efros [58] introduced the "Name That Dataset" game to discover dataset bias. Tommasi et al. [57] and Liu and He [34] expanded on this by applying convolutional neural networks and large-scale datasets, respectively. Existing work on dataset bias focuses on classification to highlight biases in data collection for building unbiased datasets, our work leverages distribution classification to explore a fundamental question: How far is the generated distribution from the real distribution? Besides, we focus on the discrepancy between the generated distribution and the real distribution rather than between specific datasets. Furthermore, the conclusion drawn by related work highlights the ability of neural networks to detect biases within increasingly general datasets, our conclusion underscores the significant gap that still exists between current diffusion models and the real data distribution. We also highlight the potential of distribution classification methods for gaining deeper insights into generative models and more effectively leveraging generated data.
Generated image detection. Several empirical works on generated image detection span three key areas: universal detection [41, 36, 40, 60, 20], reconstruction-based methods [63, 38], and frequency analysis [59, 68, 15, 6, 19, 16, 64, 11, 47, 10]. Prior research typically focuses on detecting images from various generative models trained on diverse datasets, often without complete access to the training data. Consequently, simple classification methods prove ineffective, necessitating more complex detection techniques. In contrast, our work utilizes a distinct setup where the classifier is trained with both generated and original training data, enabling effective discrimination through straightforward classification. Furthermore, our primary objective is not universal generated image detection, but rather quantifying the distributional discrepancy between real and generated data. Besides, instead of prioritizing classifier performance, we focus on the insights provided into generative model behavior and limitations.
Training with generated images. Several studies have explored the use of generated data to enhance classification performance [70, 2, 24]. These works typically leverage pre-trained text-to-image diffusion models, such as Stable Diffusion [48]. In contrast, our experiments exclusively utilize generative models trained on the original training data. Other research has investigated learning visual representations solely from synthetic images and captions [56, 18], demonstrating the viability of using synthetic data and achieving strong performance even with purely generated data. Our findings indicate that augmenting real data with generated data yields better results than complete replacement.
@@ 3 Experimental settings
We present the main experimental settings as follows. For more details, please see Appendix B.
Dataset. For label-to-image tasks, we use the CIFAR-10 [30] and ImageNet [12] datasets, and for text-to-image tasks, we utilize the COCO2014 dataset [33]. For CIFAR-10, the real distribution is constructed from its training and testing sets, comprising 50k training images and 10k testing images. The generated distribution is created using diffusion models to produce an equivalent number of images. For ImageNet, we consider resolutions of 256 and 512 and follow ADM’s [13] method to preprocess ImageNet into the desired resolutions. We randomly sample 100k training images and 50k validation images from the corresponding ImageNet sets to represent the real distribution, and generate an equivalent number of images using diffusion models for the generated distribution. For COCO2014, we randomly sample 10k training images and 1k validation images from the respective sets to form the real distribution, and use the corresponding prompts to generate an equivalent number of data for the generated distribution.
Classifier. By default, we use ResNet-50 [22] as the classifier architecture, and also consider ConvNeXt-T [35] and ViT-S [14] for completeness. Our pre-processing follows standard supervised training [35]. Specifically, during training, the classifiers process randomly augmented crops of $224\times 224$ images. For validation, images are resized to 256 pixels on the shorter side, preserving the aspect ratio, and then center cropped to $224\times 224$. For CIFAR-10 experiments, we initialize the classifier with ResNet-50 pre-trained on ImageNet and also provide the results trained from scratch (see Appendix D.2). For ImageNet, we train ResNet-50 from scratch.
Diffusion model. For CIFAR-10 generation, we consider two diffusion models: EDM [27] and U-ViT [3]. For ImageNet-256, we explore three diffusion models: EDM2 [28], U-ViT-H/2 [3], and DiT-XL/2 [42]. Since EDM2 is originally designed for ImageNet-512 generation, we resize its generated images from 512 to 256 resolution, with EDM2-XXL achieving an FID of 2.14, comparable to U-ViT-H/2 and DiT-XL/2. For ImageNet-512 generation, we use EDM2-XXL [28]. For COCO generation, we consider three text-to-image diffusion models: Pixart-$\alpha$ [8], SDXL [43], and Playground-v2.5 [32].
@@
Abbreviations and Notation. In the figures and tables following this paper, the following abbreviations are used: CIFAR-10 (C), ImageNet (I), U-ViT (U), EDM (E), DiT-XL/2 (D). Besides, "Distribution combinations" refer to multiple distributions, where the number of distributions corresponds to the number of classes in distribution classification.
@@ 4 Generated distributions are easily classified as generated
In this section, we explore how the distribution generated by diffusion models, known for their success in image generation, differs from the real distribution. To analyze this, we use a distribution classification task to assess the extent of the difference. Besides, we also extend our evaluation to other generative models, such as GANs, in Appendix D.1.
@@ 4.2 Classifiers achieve high accuracy across various settings
Label-to-image. As shown in Tab. 2, we report the binary distribution classification accuracy for various combinations of real and generated distributions in the label-to-image scenario. Surprisingly, despite the low FID scores achieved by diffusion models, neural network-based classifiers still identify significant differences between the generated and real distributions. Specifically, all classifiers easily distinguish real from generated images across different dataset combinations, with most achieving over 98% accuracy in distribution classification. Moreover, we observe a positive correlation between distribution classification accuracy and the number of training samples for the combination of U-ViT-H/2 and ImageNet-256 (see Fig. 2). This suggests that as sample size increases, classifiers become more effective at distinguishing between distributions. Notably, with only 50k samples per distribution (less than 10% of the ImageNet samples used to train the generator), classifiers achieve over 99.5% accuracy. Furthermore, to investigate the impact of data diversity on distribution classification, we scale up the sample numbers per distribution. Using U-ViT-H/2, we generate 1.28 million images, combined with the full ImageNet training set, the classifiers achieve 100% accuracy. Besides, we extend our evaluation to other generative models, including GANs and Flow Matching Models (see results in Appendix. D.1).
Distribution combinations Classifier Accuracy (%)
U [3], U ResNet-50 50.83
ViT-S 50.57
ConvNeXt-T 50.18
E [27], E ResNet-50 50.67
ViT-S 50.44
ConvNeXt-T 50.14
Table 1: Binary distribution classification on the same distribution. Classifiers are indeed unable to distinguish samples from the same distribution, demonstrating the validity of our approach.
Distribution combinations Classifier Accuracy (%)
C, U [3] ResNet-50 99.92
@@
Table 2: Binary distribution classification on label-to-image. All classifiers yield high accuracy on various combinations of datasets and generative models. FID values for each generative model are: U: 3.11, E: 1.79, U-H/2: 2.29, and E2-XXL: 1.81. Indicates experiments conducted on the full ImageNet dataset.
Text-to-image. As shown in Tab. 3, we report the distribution classification accuracy in the text-to-image scenario. We consider combinations of four distributions: COCO [33], Pixart-$\alpha$ [8], SDXL [43], and Playground-v2.5 [32]. In this task, classifiers are not only required to distinguish between real and generated distributions but also correctly identify the specific generated distribution an image comes from. Remarkably, classifiers achieve an accuracy of over 76% across all settings. Furthermore, we observe that increasing the training sample size from 5k to 10k results in significant accuracy gains, consistent with the patterns seen in the label-to-image scenario.
4.3 Self-supervised classifiers can also identify generated images
In addition to the supervised learning protocol, we also explore self-supervised learning. Specifically, we use self-supervised pre-trained models, such as MAE [23] and MoCo v3 [9], to extract image features followed by training a linear classifier on these features, referred to as the self-supervised classifier. The distribution classification accuracy of these classifiers is in Tab. 4. Notably, the self-supervised classifiers also demonstrate a strong ability to distinguish between real and generated distributions. Moreover, we observed two intriguing phenomena: First, unlike in semantic classification tasks on ImageNet, the masked image modeling-based MAE outperforms the contrastive learning-based MoCo v3 in identifying generated images related to ImageNet. Second, while self-supervised classifiers achieve relatively high accuracy on ImageNet-related tasks, their performance drops significantly when applied to CIFAR-10. We hypothesize this decline can be contributed to two factors: (1) MAE and MoCo v3 are pre-trained on ImageNet, not CIFAR-10, and (2) diffusion models perform exceptionally well on CIFAR-10 image generation, making it challenging for a simple linear classifier to differentiate between real and generated images. Besides, we also evaluate self-supervised classifiers in text-to-image scenarios (see results in Appendix. D.3).